a grocer mixed nuts worth $.80 per pound with nuts worth $.50 per pound. How many pounds of each did he use to make a mixture of 30 pounds to sell at $.75 per pound?
Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

price ---------------- quantity
Nuts type I 0.8 ---------------- x lb
Nuts Type II 0.5 ------ 30 - x lb
Mixture 0.75 ---------------- 30
30
0.8 x + 0.5 ( 30 - x ) = 0.75 * 30
80 x + 50 ( 30 - x ) = 2250
80 x + 1500 - 50 x = 2250
80 x - 50 x = 2250 - -1500
30 x = 750
/ 30
x = 25 lb 0.80 Nuts type I
5 lb 0.50 Nuts Type II

m.ananth@hotmail.ca


Answer by stanbon(75887)

a grocer mixed nuts worth $.80 per pound with nuts worth $.50 per pound. How many pounds of each did he use to make a mixture of 30 pounds to sell at $.75 per pound?
==================
Using 2 equations:
Quantity: a + b = 30 lbs
Value:::80a + 50b = 75*30
------------
Modify the equations for elimination:
8a + 8b = 8*30
8a + 5b = 3*75
------
Subtract and solve for "b":
3b = 15
b = 5 lbs (amt of 50 cent nuts)
--------
Solve for "a" using a + b = 30
a + 5 = 30
a = 25 lbs (amt. of 80 cent nuts)
==============
Cheers,
Stan H.
================